OpenStudy (anonymous):

If you bought a car at $1900 which depreciates 20% each year, what will it be worth after 4 years? A. $578 B. $778 C. $988 D. $1,588

OpenStudy (anonymous):

@ninjawolfjj

OpenStudy (anonymous):

@e.cociuba

jimthompson5910 (jim_thompson5910):

A = P(1 - r)^t A = 1900(1 - 0.2)^4 A = ???

OpenStudy (anonymous):

778.24

jimthompson5910 (jim_thompson5910):

so it's B

OpenStudy (anonymous):

How much will $600 amount to be if it is invested at 10% interest compounded semiannually (two times a year) for 4 years? A. $886.47 B. $952.40 C. $1063.20 D. $1168.17

jimthompson5910 (jim_thompson5910):

A = P(1+r/n)^(n*t) plug in P = 600 r = 0.10 n = 2 t = 4

OpenStudy (anonymous):

one sec about to solve it

jimthompson5910 (jim_thompson5910):

k

OpenStudy (anonymous):

886

OpenStudy (anonymous):

If the current value of a property is $20,000 and its value is increasing at 8% per year. Find the expression for the value after t years. A. $20,000(1.08)t B. $20,000(1.8)t C. $20,000(0.92)t D. $20,000(1.18)t

jimthompson5910 (jim_thompson5910):

886 is close, it should be 886.47, but you got the idea

jimthompson5910 (jim_thompson5910):

for this last one, you got ???

OpenStudy (anonymous):

yep I knew that I new it was a lol

OpenStudy (anonymous):

for last one I got C

jimthompson5910 (jim_thompson5910):

1+r = 1+0.08 = 1.08, so it should be A

jimthompson5910 (jim_thompson5910):

you would do 1-r = 1-0.8 = 0.92 if it were decreasing, but it's INCREASING

OpenStudy (anonymous):

okay last one: Which of the following functions increases, going upwards from left to right on a graph? A. y = 2(1/3)x B. y = -2(3)x C. y = 2x + 6 D. y = 3(1/2)x

jimthompson5910 (jim_thompson5910):

you tell me

OpenStudy (anonymous):

d

jimthompson5910 (jim_thompson5910):

nope, this is a decreasing exponential function

OpenStudy (anonymous):

B

jimthompson5910 (jim_thompson5910):

if it were positive, then it would be increasing

OpenStudy (anonymous):

yep so that would make it B, because its a negative right

jimthompson5910 (jim_thompson5910):

it can't be B because B is decreasing

jimthompson5910 (jim_thompson5910):

if B was positive (it's not), then it would be increasing

OpenStudy (anonymous):

oh okay, then it must be A

jimthompson5910 (jim_thompson5910):

A isn't it for the same reason as D

jimthompson5910 (jim_thompson5910):

so you can probably guess what it must be lol

OpenStudy (anonymous):

c lol:)

jimthompson5910 (jim_thompson5910):

yep

OpenStudy (anonymous):

@jim_thompson5910 can you help me with this assignment, im very confused.

jimthompson5910 (jim_thompson5910):

sry don't have enough time to do the whole thing

OpenStudy (anonymous):

could u possibly tomorrow:)?

jimthompson5910 (jim_thompson5910):

i might have more time tomorrow